OpenStudy (aaronandyson):

a boy walks to his school at a dist of 6 km at a constant speed of 2.5kmhr and returns home on the same path at a constant speed of 4 kmhr ...find avg vel

OpenStudy (mayankdevnani):

your work?

OpenStudy (aaronandyson):

total dist = 12 constant speed = 2.5+4 = 6.5 12/6.5 = 120/65 = 24/13

OpenStudy (mayankdevnani):

is answer 0 ?

OpenStudy (aaronandyson):

its 40/13..but how?

OpenStudy (mayankdevnani):

bhai displacement toh zero hoga na

OpenStudy (mayankdevnani):

avg velocity poochi h

OpenStudy (aaronandyson):

that i know....

OpenStudy (mayankdevnani):

phir answer 0 hona chahiye na

OpenStudy (aaronandyson):

aur avg speed kya hogi?

OpenStudy (mayankdevnani):

\[\large \bf avg~speed=\frac{total~distance}{total~time~taken}\]

OpenStudy (aaronandyson):

time nahi diya h...

OpenStudy (mayankdevnani):

time=distance/speed

OpenStudy (mayankdevnani):

\[\large \bf =\frac{6+6}{\frac{6}{2.5}+\frac{6}{4}}\]

OpenStudy (mayankdevnani):

solve kr le

OpenStudy (aaronandyson):

solve karta hon..let me frame d egn..